OpenStudy (anonymous):

how do you make x^2-17=8y+39 into vertex form?

OpenStudy (precal):

you must complete the square

OpenStudy (anonymous):

and how do you do that?

OpenStudy (anonymous):

\[y =\frac{ 1 }{ 8 }\left( x ^{2}-56 \right)\]or\[x ^{2}=8\left( y+7 \right)\]Please mark as "Best Answer" if it helps!

OpenStudy (anonymous):

but how did you get there?

OpenStudy (anonymous):

http://bit.ly/rA7QkA Ask me if you need more help.

OpenStudy (precal):